est: lest IwO ning: 00:50:45 This Question: 1 pt 39 of 47 (38 complete) This Test: 50 pts possib Cost of debt using both methods (YTM and the approximation formula) Currenty, Warren Industries can sell 15-year, $1,000-par value bonds paying annual interest at a 7% coupon rate Because current market rates or similar bonds are just under 7%, Warren can sell its bonds or $1,010 each, Warren costs of S30 per bond. The firm is in the 21% tax bracket icur no aton a. Find the net proceeds from the sale of the bond, Na b. Calculate the bonds yield to matunity (YTM) to estimate the before-tax and after-tax costs of debt c. Use the approximation formula to estimate the before-tax and after-tax costs of debt a. The net proceeds from the sale of the bond, No is $980l (Round to the nearest dollar) Round to two decimal places ) b. Using the bonds YTM, the before-tax cost of debt is Using the bonds YTM, the after tax cost of debts»% (Round to two decimal places ) c. Using the approximation Iormula, the bet re tax cost of det trs % (Round to two decimal places) Using the appro maton torm uta, moater tax cost ot debt is % (Round to two dec mal places ) Enter your answer in each of the answer boxes O Type here t Hi e
38 of 47 (38 complete) Previous Question are commonly issued in the reorganization of a failed or failing firm O A. Floating rate bonds O B. Income bonds O C. Equipment trust certificates O D. Mortgage bonds O Type here
Answer #1

1. a. Net proceed (Nd) = selling price per bond - flotation cost per bond

= $1010 - $30

= $980

b. Yield to maturity (before tax) = Coupon payment + {(net proceed - par value) / maturity} / (net proceed + par value) /2)

= (1000*7%) +  {(980- 1000) /15} / (980 + 1000 /2)

= (70- 1.3333) / 990

= 68.6667 / 990

= 6.94%

Yield to maturity (after tax) = before tax yield * (1-tax rate)

= 6.94% * (1- 0.21)

= 5.48%

2. Income Bonds ,It is that type of issued bonds in which only principal value or par value of bond will be paid to investors.Interest payment will be there if the company has enough earnings to pay so